OpenStudy (anonymous):

Solve x2 + 4x – 12 = 0 by completing the square. Show your work for full credit.

OpenStudy (anonymous):

x^2+4x-12=0 (x+6)(x-2)=0 x+6=0; x-2=0 x=-6 and x=2

OpenStudy (anonymous):